OpenStudy (anonymous):

Suppose a magnetic field of 1.0 T is applied to an area within a circular loop of wire. The strength of the field is then increased to 4.0 T. By what factor is the magnetic flux changed? A. 2 B. 4 C. 0.5 D. 0.25

OpenStudy (michele_laino):

magnetic flux is given by the subsequent formula: \[\Large \Phi = BS\]

OpenStudy (michele_laino):

where S is the area enclosed by our circular loop

OpenStudy (anonymous):

ok! what do we plug in?

OpenStudy (michele_laino):

so we have: \[\Large \begin{gathered} {\Phi _1} = 1 \times S = S \hfill \\ {\Phi _2} = 4 \times S = 4S \hfill \\ \hfill \\ \frac{{{\Phi _2}}}{{{\Phi _1}}} = \frac{{4S}}{S} = 4 \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

ohhh so our solution is 4? :O

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

yay! thanks!:)

OpenStudy (michele_laino):

:)